OpenStudy (anonymous):

How would you write a translation that maps P(-4, 2) onto the point P ' (-1, -1)? I'm confused

OpenStudy (anonymous):

the translation rule here is x'=x+3 y'=y-3

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

youw...